# Enhancing Large Language Models for Time-Series Forecasting via Vector-Injected In-Context Learning

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
As an ever-evolving complex system, the World Wide Web requires predictive capabilities to proactively address shifts in user behavior and usage patterns, thereby achieving efficient and stable operation (Kamarthi et al., 2022; Huang et al., 2025). In this process, Time-series forecasting (TSF) plays a pivotal role (Liu et al., 2024a). The goal of TSF is to forecast future trends and dynamics using historical data collected across the network (Wu et al., 2021). The accuracy of forecasting provides important support for the development of intelligent network services, encompassing core application scenarios such as personalized content recommendation (Zhang et al., 2015), network economic modeling (Barigozzi and Brownlees, 2019), and infrastructure resource allocation (Liu et al., 2022). In recent years, with the rapid development of large language models (LLMs), the research community has begun to explore applying LLMs to TSF. These approaches (referred to as LLM4TSF) (Jin et al., 2024; Liu et al., 2024b, 2025; Chang et al., 2025) have demonstrated strong predictive performance across multiple forecasting benchmarks.
@@ 3.1. Time Series Forecasting (TSF)
Here, $T_{h}$ is referred to as the look-back length, while $T_{f}$ is the prediction length.
@@ 3.2. Residual Stream
where $a_{l}^{t}$ is the output of the multi-head attention (MHA) module at layer $l$ and position $t$, and $m_{l}^{t}$ is the output of the multi-layer perceptron (MLP) at layer $l$ and position $t$. Figure 1 presents the MHA and MLP operations of a Transformer layer from the perspective of the residual stream, making it easier for readers to understand. As Figure 1 shows, $r_{l}^{t}$ is essentially the output hidden state at layer $l$ and at the $t$-th token position. The MHA and MLP can be regarded as operators that interact with the residual stream: both read from the stream and write their outputs back into it via Eq.3, thereby enabling the addition and deletion of information within the stream. Building on this perspective, many studies (Højer et al., 2025; Niu et al., 2024; Panickssery et al., 2023) directly add specific features into the residual stream to intervene in LLM behavior, steering models toward desired outcomes (e.g., “jailbreaking” or injecting specific “memory”). For more related work to the residual stream, see Appendix E. Inspired by these methods, we intervene in the LLM by adding the context vector (Eq.7) into the residual stream, thereby making it better suited for TSF.
@@ 6.1.3. Implementation Details:
All experiments are conducted on NVIDIA A100-80G GPUs using PyTorch (Imambi et al., 2021). We adopt the Adam optimizer (Kingma and Ba, 2014) with initial learning rates selected from {3e-5, 5e-5, 7e-5, 9e-5}, and use MSE loss for optimization. Our method makes predictions in an autoregressive manner and transforms the multivariate data into univariate data by treating each feature of the sequence as an individual time series. The setting of the length of the patch $P$ can be found in Table 7. The look-back length for ETT, Weather, ECL, and Traffic is set to 672. The look-back length for M3 and M4 is set to twice the forecast length. The forecast lengths for each dataset are listed in Table 6. We adopt LLaMA-7B (Touvron et al., 2023) as our base LLM, and we also conduct additional experiments with other LLMs to validate the effectiveness of our approach (Section 6.5.2). The batch size is chosen from {192, 256, 1024, 2048}, and the number of training epochs is capped at 40. Early stopping is triggered if the validation loss fails to decrease for three consecutive evaluations. For a single example $(X_{s},Y_{s})$ in Eq.4, the length of $X_{s}$ is the same as the look-back length of each dataset, and the length of $Y_{s}$ is the same as the minimum predicted length of each dataset. For each dataset, we randomly sample 5% of the training data to generate context vectors; its effectiveness is discussed in Section C.2. The context vector can be generated before training and does not affect the model’s subsequent inference speed.
@@ 6.2. Standard Dataset Evaluation
Dataset LVICL AutoTimes TimeLLM FPT Unitime iTransformer DLinear PatchTST TimesNet
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.381 0.412 0.389 0.422 0.408 0.423 0.427 0.426 0.442 0.448 0.438 0.450 0.423 0.437 0.413 0.431 0.458 0.450
ETTh2 0.326 0.376 0.352 0.395 0.354 0.393 0.353 0.391 0.356 0.399 0.382 0.414 0.431 0.446 0.330 0.379 0.414 0.427
ETTm1 0.328 0.378 0.332 0.380 0.350 0.378 0.366 0.382 0.375 0.403 0.370 0.399 0.357 0.378 0.351 0.380 0.400 0.406
ETTm2 0.239 0.306 0.243 0.310 0.254 0.314 0.265 0.315 0.277 0.329 0.272 0.331 0.267 0.333 0.255 0.315 0.291 0.333
ECL 0.158 0.248 0.159 0.253 0.159 0.253 0.167 0.263 0.216 0.305 0.161 0.256 0.177 0.274 0.159 0.253 0.192 0.295
Weather 0.224 0.262 0.235 0.273 0.225 0.257 0.237 0.270 0.253 0.276 0.238 0.272 0.240 0.300 0.226 0.264 0.259 0.287
Traffic 0.370 0.260 0.374 0.264 0.388 0.264 0.414 0.294 0.429 0.304 0.379 0.272 0.434 0.295 0.391 0.264 0.620 0.336
@@
Category Metrics LVICL AutoTimes TimeLLM FPT DLinear PatchTST TimesNet
@@
Figure 5. Impact of the injection location of the context vector.
@@
Table 5. Training time (ms/iter) with batch size fixed at 512.
Base Model Training Time (ms/iter)
GPT-2 102.8
OPT-250M 180.6
OPT-1.3B 286.1
OPT-2.7B 383.3
LLaMA-7B 758.3
@@ 6.5.2. Method generality
Additionally, with the batch size fixed at 512, we evaluate the training speed of our method when using different LLMs as the base model. The results are in Table 5. Together with Table 5 (model efficiency) and Table 4 (forecasting performance), these results allow readers to trade off predictive accuracy against efficiency.
@@ 7. Conclusion
In this paper, we propose LVICL. With all LLM parameters frozen, LVICL recovers most of the gains of full fine-tuning while markedly reducing GPU memory usage. Specifically, in LVICL, we first integrate the LLM with a learnable context-vector adapter to adaptively extract a context vector from multiple examples. We then add this vector to the residual stream of the LLM during the forward pass to improve predictive performance. Experiments show that LVICL substantially outperforms partial-parameter fine-tuning methods such as LoRA and is only slightly inferior to full fine-tuning (Figure 7); meanwhile, during training, it requires less GPU memory than both (particularly full fine-tuning). Across multiple real-world datasets, LVICL delivers competitive forecasting results.
@@ Appendix A Dataset Descriptions
  • Weather (Wu et al., 2021): Contains 21 meteorological variables sampled every 10 minutes, collected from the meteorological station of the Max Planck Institute for Biogeochemistry in 2020.
  • ECL (Wu et al., 2021): Electricity consumption data from 321 users, recorded hourly.
  • Traffic (Wu et al., 2021): Road occupancy rates collected from 862 sensors on highways in the San Francisco Bay Area, recorded hourly, covering January 2015 to December 2016.
@@
Figure 8. Performance of our method with different look-back lengths.
Figure 9. The predictive performance of our method when using different proportions of training data as examples.
@@
Dataset Prediction LVICL AutoTimes TimeLLM FPT Unitime iTransformer DLinear PatchTST TimesNet
Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.351 0.389 0.360 0.400 0.362 0.392 0.376 0.397 0.397 0.418 0.386 0.405 0.375 0.399 0.370 0.399 0.384 0.402
192 0.379 0.408 0.388 0.419 0.398 0.418 0.416 0.418 0.434 0.439 0.422 0.439 0.405 0.416 0.413 0.421 0.557 0.436
336 0.392 0.417 0.401 0.429 0.430 0.427 0.442 0.433 0.468 0.458 0.444 0.457 0.439 0.443 0.422 0.436 0.491 0.469
720 0.402 0.434 0.406 0.440 0.442 0.457 0.477 0.456 0.469 0.477 0.500 0.498 0.472 0.490 0.447 0.466 0.521 0.500
Avg 0.381 0.412 0.389 0.422 0.408 0.423 0.427 0.426 0.442 0.448 0.438 0.450 0.423 0.437 0.413 0.431 0.458 0.450
ETTh2 96 0.270 0.332 0.282 0.342 0.288 0.341 0.287 0.341 0.297 0.354 0.304 0.360 0.289 0.353 0.274 0.336 0.340 0.374
192 0.333 0.378 0.348 0.387 0.351 0.389 0.350 0.383 0.365 0.404 0.377 0.403 0.383 0.418 0.339 0.379 0.402 0.414
336 0.325 0.377 0.365 0.412 0.362 0.401 0.373 0.395 0.359 0.401 0.405 0.429 0.448 0.465 0.329 0.380 0.452 0.452
720 0.375 0.416 0.412 0.440 0.415 0.440 0.401 0.443 0.403 0.436 0.443 0.464 0.605 0.551 0.379 0.422 0.462 0.468
Avg 0.326 0.376 0.352 0.395 0.354 0.393 0.353 0.391 0.356 0.399 0.382 0.414 0.431 0.446 0.330 0.379 0.414 0.427
ETTm1 96 0.269 0.343 0.274 0.343 0.284 0.341 0.301 0.343 0.308 0.358 0.312 0.366 0.299 0.343 0.290 0.342 0.338 0.375
192 0.309 0.371 0.316 0.370 0.327 0.363 0.348 0.362 0.354 0.391 0.347 0.385 0.335 0.365 0.332 0.369 0.374 0.387
336 0.343 0.390 0.344 0.390 0.368 0.387 0.386 0.397 0.396 0.413 0.379 0.404 0.369 0.386 0.366 0.392 0.410 0.411
720 0.391 0.408 0.392 0.418 0.423 0.419 0.430 0.427 0.442 0.450 0.441 0.442 0.425 0.421 0.416 0.420 0.478 0.450
Avg 0.328 0.378 0.332 0.380 0.350 0.378 0.366 0.382 0.375 0.403 0.370 0.399 0.357 0.378 0.351 0.380 0.400 0.406
ETTm2 96 0.155 0.246 0.159 0.250 0.168 0.250 0.172 0.258 0.179 0.265 0.179 0.271 0.167 0.269 0.165 0.255 0.187 0.267
192 0.210 0.282 0.213 0.285 0.216 0.294 0.231 0.287 0.239 0.309 0.242 0.313 0.224 0.303 0.220 0.292 0.249 0.309
336 0.258 0.317 0.261 0.323 0.269 0.323 0.282 0.329 0.293 0.346 0.288 0.344 0.281 0.342 0.274 0.329 0.321 0.351
720 0.335 0.378 0.341 0.381 0.363 0.387 0.373 0.387 0.398 0.397 0.378 0.397 0.397 0.421 0.362 0.385 0.408 0.403
Avg 0.239 0.306 0.243 0.310 0.254 0.314 0.265 0.315 0.277 0.329 0.272 0.331 0.267 0.333 0.255 0.315 0.291 0.333
ECL 96 0.126 0.224 0.129 0.225 0.131 0.224 0.139 0.238 0.196 0.287 0.132 0.227 0.153 0.237 0.129 0.222 0.168 0.272
192 0.148 0.235 0.147 0.241 0.152 0.241 0.153 0.251 0.199 0.291 0.153 0.249 0.152 0.249 0.147 0.240 0.184 0.289
336 0.161 0.252 0.162 0.258 0.160 0.248 0.169 0.266 0.214 0.305 0.167 0.262 0.169 0.267 0.163 0.259 0.198 0.300
720 0.196 0.282 0.199 0.288 0.192 0.298 0.206 0.297 0.254 0.335 0.254 0.335 0.233 0.344 0.197 0.290 0.220 0.320
Avg 0.158 0.248 0.159 0.253 0.159 0.253 0.167 0.263 0.216 0.305 0.161 0.256 0.177 0.274 0.159 0.253 0.192 0.295
Weather 96 0.146 0.196 0.153 0.203 0.147 0.201 0.162 0.212 0.171 0.214 0.163 0.211 0.152 0.237 0.149 0.198 0.172 0.220
192 0.190 0.240 0.201 0.250 0.189 0.234 0.204 0.248 0.217 0.254 0.205 0.250 0.220 0.282 0.194 0.241 0.219 0.261
336 0.240 0.280 0.256 0.293 0.262 0.279 0.254 0.286 0.274 0.293 0.254 0.289 0.265 0.319 0.245 0.282 0.280 0.306
720 0.312 0.330 0.331 0.345 0.304 0.316 0.326 0.337 0.329 0.340 0.329 0.340 0.323 0.362 0.314 0.334 0.365 0.359
Avg 0.224 0.262 0.235 0.273 0.225 0.257 0.237 0.270 0.253 0.276 0.238 0.272 0.240 0.300 0.226 0.264 0.259 0.287
Traffic 96 0.338 0.245 0.343 0.248 0.362 0.248 0.388 0.282 0.401 0.291 0.351 0.257 0.410 0.282 0.360 0.249 0.593 0.321
192 0.360 0.251 0.362 0.257 0.374 0.247 0.407 0.290 0.432 0.298 0.364 0.265 0.423 0.287 0.379 0.256 0.617 0.333
336 0.382 0.259 0.379 0.266 0.385 0.271 0.412 0.294 0.436 0.305 0.382 0.273 0.436 0.296 0.392 0.264 0.629 0.333
720 0.401 0.285 0.413 0.284 0.430 0.288 0.450 0.312 0.447 0.322 0.420 0.292 0.466 0.315 0.432 0.286 0.640 0.330
Avg 0.370 0.260 0.374 0.264 0.388 0.264 0.414 0.294 0.429 0.304 0.379 0.272 0.434 0.295 0.391 0.264 0.620 0.336

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792374isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
As an ever-evolving complex system, the World Wide Web requires predictive capabilities to proactively address shifts in user behavior and usage patterns, thereby achieving efficient and stable operation (Kamarthi et al., 2022). In this process, Time-series forecasting (TSF) plays a pivotal role (Ma et al., 2025). The goal of TSF is to forecast future trends and dynamics using historical data collected across the network (Zhang et al., 2024; Huang et al., 2025). The accuracy of forecasting provides important support for the development of intelligent network services, encompassing core application scenarios such as personalized content recommendation (Zhang et al., 2015), network economic modeling (Barigozzi and Brownlees, 2019), and infrastructure resource allocation (Liu et al., 2022). In recent years, with the rapid development of large language models (LLMs), the research community has begun to explore applying LLMs to TSF. These approaches (referred to as LLM4TSF) (Jin et al., 2024; Liu et al., 2024b, 2025; Chang et al., 2025) have demonstrated strong predictive performance across multiple forecasting benchmarks.
@@ 3.1. Time Series Forecasting (TSF)
Here, $T_{h}$ is the look-back length, while $T_{f}$ is the prediction length.
@@ 3.2. Residual Stream
where $a_{l}^{t}$ is the output of the multi-head attention (MHA) module at layer $l$ and position $t$, and $m_{l}^{t}$ is the output of the multi-layer perceptron (MLP) at layer $l$ and position $t$. Figure 1 presents the MHA and MLP operations of a Transformer layer from the perspective of the residual stream, making it easier for readers to understand. As Figure 1 shows, $r_{l}^{t}$ is essentially the output hidden state at layer $l$ and at the $t$-th token position. The MHA and MLP can be regarded as operators that interact with the residual stream: both read from the stream and write their outputs back into it via Eq.3, thereby enabling the addition and deletion of information within the stream. Building on this perspective, many studies (Højer et al., 2025; Niu et al., 2024; Panickssery et al., 2023) directly add specific features into the residual stream to intervene in LLM behavior, steering models toward desired outcomes (e.g., “jailbreaking” or injecting specific “memory”). Inspired by these methods, we intervene in the LLM by adding the context vector (Eq.7) into the residual stream, thereby making it better suited for TSF.
@@ 6.1.3. Implementation Details:
All experiments are conducted on NVIDIA A100-80G GPUs using PyTorch (Imambi et al., 2021). We adopt the Adam optimizer (Kingma and Ba, 2014) with initial learning rates selected from {3e-5, 5e-5, 7e-5, 9e-5}, and use MSE loss for optimization. Our method makes predictions in an autoregressive manner and transforms the multivariate data into univariate data by treating each feature of the sequence as an individual time series. The setting of the length of the patch $P$ can be found in Table 6. The look-back length for ETT, Weather, ECL, and Traffic is set to 672. The look-back length for M3 and M4 is set to twice the forecast length. The forecast lengths for each dataset are listed in Table 5. We adopt LLaMA-7B (Touvron et al., 2023) as our base LLM, and we also conduct additional experiments with other LLMs to validate the effectiveness of our approach (Section 6.5.2). The batch size is chosen from {192, 256, 1024, 2048}, and the number of training epochs is capped at 40. Early stopping is triggered if the validation loss fails to decrease for three consecutive evaluations. For a single example $(X_{s},Y_{s})$ in Eq.4, the length of $X_{s}$ is the same as the look-back length of each dataset, and the length of $Y_{s}$ is the same as the minimum predicted length of each dataset. For Traffic and ECL, which have a large amount of training data, we randomly select 5% of the training data to generate context vectors. For other datasets, we use all the training data to generate context vectors. The context vector can be generated before training and does not affect the model’s subsequent inference speed.
@@ 6.2. Standard Dataset Evaluation
Dataset LVICL AutoTimes TimeLLM FPT Unitime iTransformer DLinear PatchTST TimesNet Chronos TimeFM
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.381 0.412 0.389 0.422 0.408 0.423 0.427 0.426 0.442 0.448 0.438 0.450 0.423 0.437 0.413 0.431 0.458 0.450 0.401 0.434 0.417 0.451
ETTh2 0.326 0.376 0.352 0.395 0.354 0.393 0.353 0.391 0.356 0.399 0.382 0.414 0.431 0.446 0.330 0.379 0.414 0.427 0.327 0.377 0.334 0.385
ETTm1 0.328 0.378 0.332 0.380 0.350 0.378 0.366 0.382 0.375 0.403 0.370 0.399 0.357 0.378 0.351 0.380 0.400 0.406 0.322 0.371 0.341 0.393
ETTm2 0.239 0.306 0.243 0.310 0.254 0.314 0.265 0.315 0.277 0.329 0.272 0.331 0.267 0.333 0.255 0.315 0.291 0.333 0.245 0.313 0.242 0.309
ECL 0.158 0.248 0.159 0.253 0.159 0.253 0.167 0.263 0.216 0.305 0.161 0.256 0.177 0.274 0.159 0.253 0.192 0.295 0.164 0.258 0.162 0.255
Weather 0.224 0.262 0.235 0.273 0.225 0.257 0.237 0.270 0.253 0.276 0.238 0.272 0.240 0.300 0.226 0.264 0.259 0.287 0.220 0.263 0.229 0.270
Traffic 0.370 0.260 0.374 0.264 0.388 0.264 0.414 0.294 0.429 0.304 0.379 0.272 0.434 0.295 0.391 0.264 0.620 0.336 0.376 0.264 0.383 0.269
@@
Scenario Category LVICL AutoTimes TimeLLM FPT DLinear PatchTST TimesNet
@@
Figure 5. Impact of context vector injection location.
@@ 7. Conclusion
In this paper, we propose LVICL. With all LLM parameters frozen, LVICL recovers most of the gains of full fine-tuning while markedly reducing GPU memory usage. Specifically, in LVICL, we first integrate the LLM with a learnable context-vector adapter to adaptively extract a context vector from multiple examples. We then add this vector to the residual stream of the LLM during the forward pass to improve predictive performance. Experiments show that LVICL significantly outperforms LoRA and approaches full fine-tuning performance with lower training memory, delivering excellent results across multiple real-world datasets.
@@ Appendix A Dataset Descriptions
  • Weather (Liu et al., 2023b): Contains 21 meteorological variables sampled every 10 minutes, collected from the meteorological station of the Max Planck Institute for Biogeochemistry in 2020.
  • ECL (Liu et al., 2023b): Electricity consumption data from 321 users, recorded hourly.
  • Traffic (Liu et al., 2023b): Road occupancy rates collected from 862 sensors on highways in the San Francisco Bay Area, recorded hourly, covering January 2015 to December 2016.
@@
Figure 8. (a) Performance of our method with different look-back lengths. (b) The predictive performance of our method when using different proportions of training data as examples.
@@
Dataset Prediction LVICL AutoTimes TimeLLM FPT Unitime iTransformer DLinear PatchTST TimesNet Chronos TimeFM
Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.351 0.389 0.360 0.400 0.362 0.392 0.376 0.397 0.397 0.418 0.386 0.405 0.375 0.399 0.370 0.399 0.384 0.402 0.369 0.409 0.384 0.426
192 0.379 0.408 0.388 0.419 0.398 0.418 0.416 0.418 0.434 0.439 0.422 0.439 0.405 0.416 0.413 0.421 0.557 0.436 0.399 0.429 0.415 0.447
336 0.392 0.417 0.401 0.429 0.430 0.427 0.442 0.433 0.468 0.458 0.444 0.457 0.439 0.443 0.422 0.436 0.491 0.469 0.413 0.439 0.429 0.456
720 0.402 0.434 0.406 0.440 0.442 0.457 0.477 0.456 0.469 0.477 0.500 0.498 0.472 0.490 0.447 0.466 0.521 0.500 0.423 0.457 0.440 0.475
Avg 0.381 0.412 0.389 0.422 0.408 0.423 0.427 0.426 0.442 0.448 0.438 0.450 0.423 0.437 0.413 0.431 0.458 0.450 0.401 0.434 0.417 0.451
ETTh2 96 0.270 0.332 0.282 0.342 0.288 0.341 0.287 0.341 0.297 0.354 0.304 0.360 0.289 0.353 0.274 0.336 0.340 0.374 0.271 0.333 0.277 0.340
192 0.333 0.378 0.348 0.387 0.351 0.389 0.350 0.383 0.365 0.404 0.377 0.403 0.383 0.418 0.339 0.379 0.402 0.414 0.326 0.371 0.334 0.380
336 0.325 0.377 0.365 0.412 0.362 0.401 0.373 0.395 0.359 0.401 0.405 0.429 0.448 0.465 0.329 0.380 0.452 0.452 0.338 0.394 0.340 0.393
720 0.375 0.416 0.412 0.440 0.415 0.440 0.401 0.443 0.403 0.436 0.443 0.464 0.605 0.551 0.379 0.422 0.462 0.468 0.372 0.409 0.384 0.426
Avg 0.326 0.376 0.352 0.395 0.354 0.393 0.353 0.391 0.356 0.399 0.382 0.414 0.431 0.446 0.330 0.379 0.414 0.427 0.327 0.377 0.334 0.385
ETTm1 96 0.269 0.343 0.274 0.343 0.284 0.341 0.301 0.343 0.308 0.358 0.312 0.366 0.299 0.343 0.290 0.342 0.338 0.375 0.264 0.337 0.280 0.357
192 0.309 0.371 0.316 0.370 0.327 0.363 0.348 0.362 0.354 0.391 0.347 0.385 0.335 0.365 0.332 0.369 0.374 0.387 0.303 0.364 0.321 0.386
336 0.343 0.390 0.344 0.390 0.368 0.387 0.386 0.397 0.396 0.413 0.379 0.404 0.369 0.386 0.366 0.392 0.410 0.411 0.337 0.383 0.357 0.405
720 0.391 0.408 0.392 0.418 0.423 0.419 0.430 0.427 0.442 0.450 0.441 0.442 0.425 0.421 0.416 0.420 0.478 0.450 0.384 0.401 0.406 0.424
Avg 0.328 0.378 0.332 0.380 0.350 0.378 0.366 0.382 0.375 0.403 0.370 0.399 0.357 0.378 0.351 0.380 0.400 0.406 0.322 0.371 0.341 0.393
ETTm2 96 0.155 0.246 0.159 0.250 0.168 0.250 0.172 0.258 0.179 0.265 0.179 0.271 0.167 0.269 0.165 0.255 0.187 0.267 0.159 0.252 0.157 0.249
192 0.210 0.282 0.213 0.285 0.216 0.294 0.231 0.287 0.239 0.309 0.242 0.313 0.224 0.303 0.220 0.292 0.249 0.309 0.215 0.288 0.212 0.285
336 0.258 0.317 0.261 0.323 0.269 0.323 0.282 0.329 0.293 0.346 0.288 0.344 0.281 0.342 0.274 0.329 0.321 0.351 0.264 0.324 0.261 0.320
720 0.335 0.378 0.341 0.381 0.363 0.387 0.373 0.387 0.398 0.397 0.378 0.397 0.397 0.421 0.362 0.385 0.408 0.403 0.343 0.387 0.338 0.382
Avg 0.239 0.306 0.243 0.310 0.254 0.314 0.265 0.315 0.277 0.329 0.272 0.331 0.267 0.333 0.255 0.315 0.291 0.333 0.245 0.313 0.242 0.309
ECL 96 0.126 0.224 0.129 0.225 0.131 0.224 0.139 0.238 0.196 0.287 0.132 0.227 0.153 0.237 0.129 0.222 0.168 0.272 0.131 0.233 0.129 0.230
192 0.148 0.235 0.147 0.241 0.152 0.241 0.153 0.251 0.199 0.291 0.153 0.249 0.152 0.249 0.147 0.240 0.184 0.289 0.154 0.244 0.152 0.241
336 0.161 0.252 0.162 0.258 0.160 0.248 0.169 0.266 0.214 0.305 0.167 0.262 0.169 0.267 0.163 0.259 0.198 0.300 0.167 0.262 0.165 0.259
720 0.196 0.282 0.199 0.288 0.192 0.298 0.206 0.297 0.254 0.335 0.254 0.335 0.233 0.344 0.197 0.290 0.220 0.320 0.204 0.293 0.201 0.290
Avg 0.158 0.248 0.159 0.253 0.159 0.253 0.167 0.263 0.216 0.305 0.161 0.256 0.177 0.274 0.159 0.253 0.192 0.295 0.164 0.258 0.162 0.255
Weather 96 0.146 0.196 0.153 0.203 0.147 0.201 0.162 0.212 0.171 0.214 0.163 0.211 0.152 0.237 0.149 0.198 0.172 0.220 0.145 0.198 0.151 0.202
192 0.190 0.240 0.201 0.250 0.189 0.234 0.204 0.248 0.217 0.254 0.205 0.250 0.220 0.282 0.194 0.241 0.219 0.261 0.188 0.242 0.196 0.248
336 0.240 0.280 0.256 0.293 0.262 0.279 0.254 0.286 0.274 0.293 0.254 0.289 0.265 0.319 0.245 0.282 0.280 0.306 0.238 0.281 0.248 0.289
720 0.312 0.330 0.331 0.345 0.304 0.316 0.326 0.337 0.329 0.340 0.329 0.340 0.323 0.362 0.314 0.334 0.365 0.359 0.309 0.331 0.322 0.340
Avg 0.224 0.262 0.235 0.273 0.225 0.257 0.237 0.270 0.253 0.276 0.238 0.272 0.240 0.300 0.226 0.264 0.259 0.287 0.220 0.263 0.229 0.270
Traffic 96 0.338 0.245 0.343 0.248 0.362 0.248 0.388 0.282 0.401 0.291 0.351 0.257 0.410 0.282 0.360 0.249 0.593 0.321 0.343 0.249 0.350 0.253
192 0.360 0.251 0.362 0.257 0.374 0.247 0.407 0.290 0.432 0.298 0.364 0.265 0.423 0.287 0.379 0.256 0.617 0.333 0.366 0.255 0.372 0.260
336 0.382 0.259 0.379 0.266 0.385 0.271 0.412 0.294 0.436 0.305 0.382 0.273 0.436 0.296 0.392 0.264 0.629 0.333 0.388 0.263 0.395 0.268
720 0.401 0.285 0.413 0.284 0.430 0.288 0.450 0.312 0.447 0.322 0.420 0.292 0.466 0.315 0.432 0.286 0.640 0.330 0.407 0.289 0.415 0.295
Avg 0.370 0.260 0.374 0.264 0.388 0.264 0.414 0.294 0.429 0.304 0.379 0.272 0.434 0.295 0.391 0.264 0.620 0.336 0.376 0.264 0.383 0.269